A frictionless cart is given a quick push that sends it rolling up a 17º incline, as shown. After the push, the cart travels 1.30 m before stopping and rolling back down the incline. How fast was the cart moving immediately after it was pushed?
Answer #1

here,

theta = 17 degree

distance travelled , s = 1.3 m

accelration , a = - g * sin(theta)

a = - 9.8 * sin(17)

a = - 2.87 m/s^2

let the initial speed be v

using third equation of motion

v^2 - u^2 = 2 * a *s

0 - u^2 = - 2 * 2.87 * 1.3

u = 2.73 m/s

the initial speed of cart after the push is 2.73 m/s